OpenStudy (anonymous):

Write an equation to describe the variation. Use k for the constant of proportionality. w varies jointly as x and the square of y.

OpenStudy (saifoo.khan):

\[w = k w y^2\]

OpenStudy (anonymous):

that's not an option for an answer :[ a. wxy^2 b. w+x+y^2=k c. w=kxy^2 d. w=k+x+y^2

OpenStudy (saifoo.khan):

that is an option coco. Check out the C part.

OpenStudy (anonymous):

you wrote w=kWy^2 C says w=kXy^2

OpenStudy (anonymous):

just a typo?

OpenStudy (saifoo.khan):

Oh! Damn, you are right. it was a typo!

OpenStudy (saifoo.khan):

\[\Huge w = k x y^2\]

OpenStudy (anonymous):

hahaha why is w by itself?

OpenStudy (saifoo.khan):

it is w on the left side coz it says, W VARIES directly.. and stuff.